given f(x)=-x-4, find f(a-3).
A). a-8
B). -a-1
C). a-1
D). -a-7

Answer by Nate(3500)

f(x) = -x - 4
f(a - 3) = -(a - 3) - 4
f(a - 3) = -a + 3 - 4
f(a - 3) = -a - 1
(B)